### v3

Abstract
The weather forecasting system is important for science and society, and significant achievements have been made in applying artificial intelligence (AI) to medium-range weather forecasting. However, existing AI-based weather forecasting models rely on analysis or reanalysis products from traditional numerical weather prediction (NWP) systems as initial conditions for making predictions. Initial states are typically generated by traditional data assimilation components, which are computational expensive and time-consuming. Here we present an AI-based data assimilation model, i.e., Adas, for global weather variables. By introducing the confidence matrix, Adas employs gated convolution to handle sparse observations and gated cross-attention for capturing the interactions between the background and observations. Further, we combine Adas with the advanced AI-based forecasting model (i.e., FengWu) to construct the first end-to-end AI-based global weather forecasting system: FengWu-Adas. We demonstrate that Adas can assimilate global observations to produce high-quality analysis, enabling the system operate stably for long term. Moreover, we are the first to apply the methods to real-world scenarios, which is more challenging and has considerable practical application potential. We have also achieved the forecasts based on the analyses generated by AI with a skillful forecast lead time exceeding that of the IFS for the first time.
Driven by the advancements and maturity of AI, particularly deep learning techniques, scientific intelligence has been rapidly evolving with the aim of leveraging AI to promote scientific research and discovery. Within the field of atmospheric science, AI has achieved remarkable achievements in various areas, such as post-processing and bias correction [1, 2, 3, 4], downscaling [5, 6], precipitation nowcasting [7, 8, 9], climate forecasting [10], and medium-range weather forecasting [11, 12, 13, 14, 15]. Some AI-based models have demonstrated highly competitive forecasts compared to the deterministic forecasts of the state-of-the-art NWP system, i.e., the Integrated Forecasting System (IFS) from the European Centre for Medium-Range Weather Forecasts (ECMWF). These models are usually trained on reanalysis dataset (e.g., ERA5 [16] from ECMWF) and allow much lower computational costs and easier deployment for operational forecasting. Despite drawbacks such as forecast smoothness and bias drift [17], AI approaches have shown the immense potential of data-driven modeling in weather prediction, offering a new paradigm for meteorological forecasting.
However, the AI-based weather forecasting models still require analysis products generated through the process of data assimilation in the traditional NWP system for making predictions (Fig. 1). Specifically, data assimilation aims to obtain the best estimate of the true state of the Earth system (known as the analysis) and provide an accurate initial state for weather prediction, thereby improving the forecast performance. With periodic corrections applied to the forecasts, data assimilation allows for a more stable and accurate long-term prediction, which is a critical component in global weather forecasting system. This process relies heavily on observations as they provide crucial information that closely represents the true state of the atmosphere. For example, the earliest initial conditions were obtained by interpolating observations onto the grid points of the grid space [18]. Modern data assimilation techniques are usually achieved by integrating observations with the short-range forecasts (i.e., the background), primarily including two categories: Kalman filters and variational methods [19]. The main motivation for introducing the background, or the first guess, lies in the incomplete nature of observations. The sparsity of observations makes it difficult to obtain a global estimate, resulting in low-quality interpolation analysis. Besides, not all state variables can be observed directly.
Currently, AI methods have primarily been applied to specific processes within data assimilation to address the limitations of traditional data assimilation techniques [20, 21]. For instance, implicit neural representations (INRs) [22] and various autoencoders (AEs) [23, 24, 25] have been employed as reduced-order-models (ROMs) to tackle high-dimensional challenges. These methods replace classical linear models such as Proper Orthogonal Decomposition (POD) [26, 27] in latent assimilation, providing an efficient framework for the representation and reconstruction of state variables in the latent space. Neural networks have also been utilized to derive tangent-linear and adjoint models in 4D variational (4D-Var) data assimilation [28], provide localization functions for ensemble Kalman filters (EnKF) [29] and estimate error covariance matrix [30, 31]. Moreover, there exist strong mathematical similarities between machine learning and data assimilation, particularly the variational data assimilation [32], which enables the optimization of the cost function in variational data assimilation using auto-differentiation [24, 33]. However, these studies have not involved altering the algorithms of data assimilation fundamentally. While traditional data assimilation algorithms are underpinned by rigorous mathematical theory and physical priors, they inevitably rely on certain assumptions as prerequisites. In practical operational applications, they often face constraints due to high computational costs, necessitating various approximations for solution [34]. In situations where supervised information is available, neural networks are expected to capture the correlations among data during the assimilation automatically and provide a new kind of algorithms [35]. Efforts have been made in some simple scenarios [36, 37, 38], but they simply feed the data into networks without any specialized design. And so far, there have been no attempts in data assimilation for real observational data in multivariate global weather scenarios.
In this study, we present Adas, a novel data assimilation model for global weather variables, which can assimilate sparse observations with different qualities and provide initial conditions for weather forecasting. Inspired by the gating mask in gated convolution [39] and the error covariance matrix in traditional data assimilation techniques, we introduce the confidence matrix to characterize the availability and quality of observations during the data assimilation. With the guidance of confidence matrix, Adas employs the gated convolution module to handle sparse observations and the gated cross-attention module for capturing the interactions between the background and observations efficiently. The prediction of the advanced AI-based weather forecasting model, FengWu [14], is used to generate the background for data assimilation. In return, the analysis of Adas is then used as the initial state of FengWu for making predictions at the next time step, thus forming the first end-to-end AI-based global weather forecasting system: FengWu-Adas. Fig. 1 shows the progression of global weather forecasting system and the main motivation of this research. We demonstrate that our system can operate stably for long term, and our method is robust and flexible to be combined with any state-of-the-art weather forecasting model without retraining or fine-tuning. The simulation experiments have proven the superior performance of FengWu-Adas in capturing the distribution patterns of the background error and producing high-quality analysis. And we further apply FengWu-Adas to real-world scenarios, which is more challenging and has considerable practical application potential. We have also achieved the forecasts based on the analyses generated by AI with a skillful forecast lead time of more than 8.75 days, which exceeds that of the IFS for the first time.
Results
Principle of FengWu-Adas
The framework of FengWu-Adas is shown schematically in Fig. 2a. For the given initial state denoted as $\dot{x}_{t}$ at time $t$, FengWu can make multi-step predictions $\hat{x}_{t+\Delta t}$, $\hat{x}_{t+2\Delta t}$, …, $\hat{x}_{t+k\Delta t}$ at time $t+\Delta t$, $t+2\Delta t$, …, $t+k\Delta t$ in an auto-regressive manner, where $k$ denotes the time index and $\Delta t=6h$ is the temporal spacing of single-step prediction. Considering that the accuracy of the background will directly affect the quality of the analysis, the single-step prediction of FengWu is used as the background $x_{t+\Delta t}^{b}$ for data assimilation. After obtaining the observations $y_{t+\Delta t}$ at the corresponding time, confidence levels ranging from 0 to 1 can be assigned to each datum based on the availability and quality of the observation, thereby generating the confidence matrix. The confidence matrix $m$, serving as supplementary information, is sent into Adas along with observations for assimilation. Subsequently, the analysis $x_{t+\Delta t}^{a}$ produced by Adas will be the initial state for the next prediction, i.e., the input of FengWu at next time and so on, in a cyclic manner. The multi-step forecasts are predicted through auto-regression based on the analysis.
Adas utilizes a dual encoder composed of gated interaction blocks to extract features of the background and sparse observations separately and capture the interactions between them (Fig. 2b). All the inputs with a shape of $H_{0}\times W_{0}\times C_{0}$ are converted into a unified latent shape of $D\times H\times W\times C$ through the patch embedding module so that the network can capture their spatial relationships in the 3D mesh. $H_{0}$ and $W_{0}$ represent the number of grid points in the longitudinal and latitudinal directions, respectively, and the channel $C_{0}$ represents the number of variables. The 3D upper-air variables and 2D surface variables are embedded into the latent space independently and then concatenated together, where the information at each grid point is represented by a $C$-dimensional vector. $H$ and $W$ correspond to the number of grid points in the longitudinal and latitudinal directions after embedding, and $D$ represents the number of levels in the vertical direction. By performing down-sampling and up-sampling through the patch merging and patch expanding module (Fig. 2d), the encoder can capture multi-scale meteorological features. The dual gated interaction block extracts features from the background and observations with standard convolution and gated convolution, along with efficient information interactions through the gated cross-attention module (Fig. 2c). Both the gated convolution and gated cross-attention are guided by the confidence matrix, which is updated together with the features to represent the availability and quality of observations (details in Methods). After the feature fusion block implemented by a convolutional layer, the features are sent into a series of Transformer [40, 41] blocks and then recovered to the original size. The convolutional operation can bring inductive bias of locality, and the attention mechanism enables the model to capture long-range dependencies. The attention block utilizes the shifted-window mechaism [42] adjusted according to geographical rules to reduce the computational cost and ensure continuity in the latitudinal direction, and $N=8$ in our implementation.
Superior performance of FengWu-Adas in simulation experiments
For testing the performance of FengWu-Adas, the Observing System Simulation Experiments (OSSEs) are performed by utilizing the ERA5 dataset firstly. The experiment proceeds with alternating weather forecasting and data assimilation steps: the single-step forecast provides the background for data assimilation, and the analysis is used as the initial state for the next forecasting step. ERA5 is regarded as the ground truth, and the simulated observations are obtained by adding a mask to the ground truth. The mask is used as the simulated confidence matrix for data assimilation. Therefore, the confidence matrix in simulation experiments is binary, with 1 indicating that the observation in corresponding grid point is available and accurate, and 0 indicating unavailable observation. The confidence matrix in real scenarios will be more complex and the confidence value is produced based on the quality of the observations. To improve the robustness of observation locations, the simulated masks for training are generated randomly with a fixed mask ratio at each time step. However, most observation locations are generally fixed in real-world scenarios, and the same strategy for testing not only deviates from reality but also poses potential issues in the cyclic experiments. Specifically, a random mask implies that any position has the probability of being observed. When the number of steps in the cyclic experiments is sufficiently large, the observation positions would approximately cover the whole world, which is clearly unrealistic. To address this concern, the masks for testing are fixed even though we find that this has almost no impact on performance in our experiments (see the comparison between Fig. 3 and Extended Data Fig. 7).
Here, we conduct three sets of experiments with different observation ratios, i.e., 90%, 99% and 99.9%. In other words, the ratios of available observations to the total number of grid points are 10%, 1% and 0.1%. The system is initialized using the ERA5 data at 2018-01-01 00:00UTC, which serves as the initial state for the first-step forecasting. Once the system is initialized, ERA5 data is invisible throughout the whole year, except for generating simulated observations. Fig. 3 illustrates the variations in RMSE and Bias (defined in Methods) of the analyses for main variables over the year. Both RMSE and Bias are computed against ERA5. The results with different observation ratios are represented by different colors, as marked in the bottom. To enhance clarity, the original data is presented with reduced opacity, while the solid lines indicate the smoothed values after the exponential moving average (EMA) with a smoothing factor of 90%. The RMSE of the analyses for all variables with all observation ratios is maintained at a low level for the whole year, showing the long-term stability of the system and the ability to generate accurate analysis based on Adas. The dashed lines are the RMSE of 6-hour forecast of the IFS, which represent fairly advanced error levels. Even when the observation ratio is only 0.1% (corresponding to an absolute number of observations of less than 0.1 million), the RMSE for all variables is lower than the dashed lines at almost all times. As the observation ratio grows, the RMSE of the analysis decreases gradually, and the Bias remains at around 0 more stably, revealing the system’s potential to eliminate bias drift of AI-based forecasting models and maintain conservation. The average RMSE over the whole year for all variables is shown in Extended Data Fig. 8.
The visualizations for z500 during the data assimilation, along with the error distributions and analysis increments, are presented in Fig. 4. The visualization date-time is randomly selected at 2018-01-26 06:00UTC, where the background is obtained from FengWu by performing single-step forecasting of the analysis at 2018-01-26 00:00UTC. The first row displays the visualizations of ERA5 (i.e., the ground truth, Fig. 4a) and the background (Fig. 4b), and the error distributions of the background (i.e., the background minus ERA5, Fig. 4c). The other rows show the visualizations of random simulated observations (the first column), the analysis increments (i.e., the analysis minus the background, the second column), and the error distributions of the analysis after data assimilation (i.e., the analysis minus ERA5, the last column) with different observation ratios. All the analysis increments exhibit a distribution similar to the background errors, indicating that Adas captures the error distribution of the background effectively. And all the errors of the analysis are reduced to varying degrees after data assimilation. As expected, the analysis increment is closer to the background error and the analysis error is smaller with the increase of observation information.
In order to further demonstrate the stability of the system, a completely random initial state is used to start the system. Extended Data Fig. 9 shows the corresponding RMSE, and the subgraph provides a clearer display of the first ten days. The y-axis in these graphs uses a coordinate axis with non equidistant logarithmic form to ensure visual perception. The first-step forecast after random initialization will produce a terrible performance, but all the errors converge quickly to a stable state with the same levels of metrics as that of starting the system with ERA5 in a few days. And the convergence speed accelerates as the number of observations increases, which is reasonable. This indicates that Adas can perform data assimilation on the background with any quality, ensuring that the system can correct back to the stable state in a short period of time without affecting subsequent operations even if it encounters a malfunction that leads to complete derailment of the forecast.
As an end-to-end global weather forecasting system that can operate independently based on observational data, FengWu-Adas can generate multi-step forecasts on its own analysis (Fig. 2a). The average RMSE and ACC skill for weather forecasting of FengWu-Adas over ten days is shown in Extended Data Fig. 10. It should be emphasized that FengWu is trained on ERA5 data [14] and is not fine-tuned to our task. Therefore, our method is flexible to be combined with any state-of-the-art weather forecasting model. Actually, our pre-trained model can assimilate observations on arbitrary background without retraining or fine-tuning, as shown in Extended Data Fig. 9. The universal capability allows Adas to be directly integrated with other weather forecasting models as a plug-and-play component, and the forecasting skill of the system will improve as the forecasting model develops.
Applying FengWu-Adas to real-world scenarios
The data assimilation for real observational data is more challenging than simulation experiments and has considerable practical application potential. To validate the performance in real-world scenarios, we apply FengWu-Adas to the assimilation for observations of Global Data Assimilation System (GDAS). The values of state variables are interpolated onto the grid points to generate the observations, and the confidence matrix is produced according to the availability and quality of the observations. Observing that satellite-derived data often have larger systematic errors, we conducted two sets of experiments to assimilate all data and non-satellite data of GDAS. Fig. 5 illustrates the variations in RMSE and Bias of the analyses for main variables over the year, where the black crosses indicate that the observational data at the corresponding time are missing. When FengWu-Adas is applied to the assimilation for real observational data, the system can still operate stably for long term, and the RMSE of some variables such as specific humidity can reach a level comparable to that of simulation experiments and the IFS (Extended Data Fig. 11). The results show that although satellite-derived data are less accurate, they can provide more information to improve the quality of the analysis, and it is expected that neural networks are robust to noise. This holds true for most variables except for the pressure level of 50hPa, which may be caused by the small number of high-altitude observational data. What’s more, Adas is robust to handle the lack of observations at a certain time. When the input observations is an all-zeros matrix, it can perform a function like post-processing to maintain the operation and produce the analysis and forecasts.
The average RMSE skill for weather forecasting based on the analyses over ten days is shown in Fig. 6, and the intersections with dashed line correspond to the skillful forecast lead times with ACC=0.6 for z500. The analyses generated after assimilating all data and non-satellite data are used as the initial states for weather forecasting, and the skillful forecast lead times reach 8.75 days and 8.25 days, respectively. The results reflect the forecasting skill of our system after assimilating real observational data, and we have achieved the forecasts based on the analyses generated by AI with a skillful forecast lead time exceeding that of the IFS (8.5 days [43]) for the first time. The GDAS experiments demonstrate the stability and considerable potential for practical applications of our system.
Discussion
There are two main reasons that make assimilating real observational data more challenging than simulated observations. On one hand, the distribution patterns of observational data are different between these scenarios. In simulation experiments, observations generated through random sampling or other methods are approximately distributed uniformly, which is an ideal distribution pattern that yields favorable performance. However, observational data often exhibit a severely non-uniform distribution due to various geographical and human factors in real-world scenarios. For instance, there is a scarcity of observations in polar regions in most cases, which can significantly impact the regional analysis. Such impacts may be propagated through the system’s operation, potentially leading to error accumulation or even collapse of the system. On the other hand, there exists an inherent inconsistency between observational values in the observational space and grid-point values in the grid space. Observational values represent the true and objective reflection of individual states at specific points in space, while grid-point values represent the overall states of a region. Therefore, observational values tend to exhibit larger fluctuations, while grid-point values are more stable and smooth theoretically. This inconsistency can be alleviated as the resolution of grid space increases and the quantity of observational data grows, but it remains a problem that cannot be ignored at present. Traditional data assimilation algorithms address this inconsistency by aligning the grid space to the observational space using observation operator. In contrast, our method requires an operator for the inverse mapping, which may be achieved by adding a pre-processing module instead of simple interpolation and further enhance the performance of our method. In summary, there exists a natural gap between simulated experiments and real-world systems. While simulation experiments are necessary for academic research and algorithm improvement, the ability to be extended to real-world scenarios is of paramount importance for practical applications. We have achieved stable long-term assimilation of real observational data for the first time, which can help forecasting models get rid of the dilemma of relying on the initial states provided by inefficient traditional data assimilation components, and promote the rapid development of research in related fields.
The future work will mainly focus on a more universal and high-performance framework. For example, techniques based on continuous space modeling [44] may provide a train of thought for better solution, which can work directly with the off-grid real observations without pre-processing or observation operator that may introduce additional errors. Nowadays, high-quality data has always been essential for the development of AI and many other technologies, and the reanalysis datasets still hold an irreplaceable significance. Therefore, developing physical models and traditional algorithms remains crucial. The purpose of our study is not to replace the traditional NWP systems but to demonstrate AI methods’ potential in tackling real-world challenges. Both physical techniques and AI methods have their own strengths and limitations, and they should be treated equally and considered as powerful tools for benefiting science and society. We believe that this is a meaningful exploration and look forward to the deployment and implementation of AI methods in operational systems in the future.
Methods
Data preparation
ERA5 is used as the ground truth and source of the simulated observations in our experiments. ERA5 is a global atmospheric reanalysis archive containing hourly weather variables such as temperature, geopotential, wind speed, humidity, etc. A subset of ECMWF’s ERA5 dataset for 40 years, from 1979 to 2018, is chosen to train and evaluate the model. We choose to conduct experiments on a total of 69 variables at a resolution of 0.25° ($721\times 1440$ grid points), including five upper-air variables with 13 pressure levels (i.e., 50hPa, 100hPa, 150hPa, 200hPa, 250hPa, 300hPa, 400hPa, 500hPa, 600hPa, 700hPa, 850hPa, 925hPa, and 1000hPa), and four surface variables. Specifically, the upper-air variables are geopotential (z), temperature (t), specific humidity (q), zonal component of wind (u) and meridional component of wind (v), whose 13 sub-variables at different vertical level are presented by abbreviating their short name and pressure levels (e.g., z500 denotes the geopotential at a pressure level of 500 hPa), and the surface variables are 10-meter zonal component of wind (u10), 10-meter meridional component of wind (v10), 2-meter temperature (t2m) and mean sea level pressure (msl). Following a common protocol in the simulation experiments, the data from 1979-2015 are used for training, 2016-2017 for validation, and 2018 for testing.
The real observational data is parsed from the prepbufr files of GDAS, which is archived by National Centers for Environmental Information (NCEI). GDAS provides various types of observational data starting from 2012-02-13 00:00UTC, including multi-source data such as surface land reports, surface marine reports, satellite-derived wind reports, etc. Each datum contains the longitude and latitude coordinates, pressure level, time, and the value and quality marker of state variables. Regarding the large number of missing GDAS observations after 2018, we use the data from 2017 for testing and others for training to demonstrate the quality of the analysis for the entire year.
Evaluation metrics
The latitude-weighted root mean square error (RMSE) and the latitude-weighted anomaly correlation coefficient (ACC) are statistical metrics widely used in geospatial analysis and atmospheric science, which are both computed against ground truth data. And latitude weighting is a common strategy to account for the varying area represented by different latitudes on a spherical Earth. RMSE evaluates the accuracy of a model’s forecasts or estimates of variables across different latitudes, and ACC reflects the model’s ability to capture anomalies (departures from the long-term averaged climatology). Regarding the bias drift of AI models, the latitude-weighted mean error (Bias) is also used to assess the bias of estimates against the ground truth [17], and the Bias of an ideal system should be close to zero. The definitions of the metrics are as below.
RMSE
Given the estimate $\hat{x}_{h,w,c}^{t}$ and its ground truth $x_{h,w,c}^{t}$ for the $c$-th channel at time $t$, the RMSE is defined as
$\operatorname{RMSE}(c,t)=\sqrt{\frac{1}{H_{0}\cdot W_{0}}\sum\nolimits_{h,w}H_{0}\cdot\frac{\operatorname{cos}(\alpha_{h,w})}{\sum_{h^{\prime}=1}^{H_{0}}\operatorname{cos}(\alpha_{h^{\prime},w})}(x_{h,w,c}^{t}-\hat{x}_{h,w,c}^{t})^{2}}$
where $h$ and $w$ denote the indices for each grid point along the longitudinal and latitudinal directions, respectively, and $\alpha_{h,w}$ is the latitude of point $(h,w)$. For the forecast $\hat{x}_{h,w,c}^{i+\tau}$ and its ground truth $x_{h,w,c}^{i+\tau}$ for the $c$-th channel with lead time $\tau$, the RMSE is defined as
$\operatorname{RMSE}(c,\tau)=\frac{1}{T}\sum_{i=1}^{T}\sqrt{\frac{1}{H_{0}\cdot W_{0}}\sum\nolimits_{h,w}H_{0}\cdot\frac{\operatorname{cos}(\alpha_{h,w})}{\sum_{h^{\prime}=1}^{H_{0}}\operatorname{cos}(\alpha_{h^{\prime},w})}(x_{h,w,c}^{i+\tau}-\hat{x}_{h,w,c}^{i+\tau})^{2}}$
where $T$ is the total number of test time slots.
ACC
Given the forecast $\hat{x}_{h,w,c}^{i+\tau}$ and its ground truth $x_{h,w,c}^{i+\tau}$, and the climatological mean over the day-of-year $C_{h,w,c}^{i+\tau}$ for the $c$-th channel with lead time $\tau$, the ACC is defined as
$\operatorname{ACC}(c,\tau)=\frac{1}{T}\sum_{i=1}^{T}\frac{\sum\nolimits_{h,w}H_{0}\cdot\frac{\operatorname{cos}(\alpha_{h,w})}{\sum_{h^{\prime}=1}^{H_{0}}\operatorname{cos}(\alpha_{h^{\prime},w})}(x_{h,w,c}^{i+\tau}-C_{h,w,c}^{i+\tau})(\hat{x}_{h,w,c}^{i+\tau}-C_{h,w,c}^{i+\tau})}{\sqrt{\sum\nolimits_{h,w}H_{0}\cdot\frac{\operatorname{cos}(\alpha_{h,w})}{\sum_{h^{\prime}=1}^{H_{0}}\operatorname{cos}(\alpha_{h^{\prime},w})}(x_{h,w,c}^{i+\tau}-C_{h,w,c}^{i+\tau})^{2}\sum\nolimits_{h,w}H_{0}\cdot\frac{\operatorname{cos}(\alpha_{h,w})}{\sum_{h^{\prime}=1}^{H_{0}}\operatorname{cos}(\alpha_{h^{\prime},w})}(\hat{x}_{h,w,c}^{i+\tau}-C_{h,w,c}^{i+\tau})^{2}}}$
where $C_{h,w,c}^{i+\tau}$ is averaged from 1993 to 2016 with ERA5 data [13, 14].
Bias
Given the estimate $\hat{x}_{h,w,c}^{t}$ and its ground truth $x_{h,w,c}^{t}$ for the $c$-th channel at time $t$, the Bias is defined as
$\operatorname{Bias}(c,t)=\frac{1}{H_{0}\cdot W_{0}}\sum\nolimits_{h,w}H_{0}\cdot\frac{\operatorname{cos}(\alpha_{h,w})}{\sum_{h^{\prime}=1}^{H_{0}}\operatorname{cos}(\alpha_{h^{\prime},w})}(x_{h,w,c}^{t}-\hat{x}_{h,w,c}^{t})$
Deep network modules
Patch embedding and patch recovery
The patch embedding and patch recovery are implemented through standard convolution and transposed convolution, respectively. The patch size of Adas in our experiments is set to 6 and the number of vertical layers remain unchanged. This means a patch has $1\times 6\times 6$ pixels for 3D upper-air variables and $6\times 6$ pixels for 2D surface variables. The kernel sizes and strides of 3D and 2D convolution are the same as the patch size, and the two parts are then concatenated together in the vertical direction. Therefore, in our experiments, $H=H_{0}/6$, $W=W_{0}/6$, $D=14$, and the dimension $C$ is 192. The parameters of patch embedding and patch recovery are not shared for each input.
Patch merging and patch expanding
In order to capture multi-scale meteorological features, the patch merging and patch expanding modules are used for down-sampling and up-sampling (Fig. 2d). The patch merging firstly rearrange pixels to reduce the horizontal dimensions by half, while the number of channels increases. After the regularization layer, the number of channels is then transformed by a linear layer. Therefore, the patch merging achieves halving of the horizontal dimensions and doubling of the channel dimension. The patch expanding module performs the opposite operation and makes the encoder form a symmetrical Unet [45] structure. Similarly, the number of vertical layers is not changed in the process of patch merging and patch expanding.
Gated convolution
Gated convolution is a convoluional operation based on soft-gating, which employs continuous masks ranging between 0 and 1 to represent the degree of validity for each pixel. Originally, the gating mask in gated convolution was learned automatically from data at each layer, and the output features were weighted by it. In Adas, the confidence matrix of our method can serve as the gating mask. By using the confidence matrix as input for each layer and updating it explicitly, we propose an improved form of gated convolution to provide a more complete representation for sparse observations. The guidance of the confidence matrix can help the network extract features more effectively and capture the interactions between the background and observations more efficiently in the gated cross-attention module. Specifically, each value in the confidence matrix signifies the degree of confidence associated with corresponding observation values, and the confidence matrix is dynamically updated to guide the varying confidence levels of different layers. Let the subscript $i$ and $o$ represent the input and output of each layer, respectively, then the gated convolution and the update rule of the confidence matrix can be formulated as
$y_{o}=SiLU\big{(}Conv3d(y_{i})\big{)}\odot Sigmoid\big{(}Conv3d(m_{i})\big{)}$ (5)
$m_{o}=Sigmoid\big{(}Conv3d(m_{i})\big{)}$ (6)
where $\odot$ represents element-wise multiplication. The $Sigmoid(\cdot)$ activation ensures that the values of the confidence matrix are between 0 and 1, and the activation funtion for features can be arbitrary. We use the $SiLU(\cdot)$ activation here, which has been proven to be a better choice than $ReLU(\cdot)$ [46, 47]. The gated convolution module can provide a more complete representation for observations and improve the quality of feature fusion.
Gated cross-attention
The cross-attention mechanism is introduced to capture the interactions between the background and observations, but treating observations at different locations with different confidence levels equally is clearly unreasonable. Especially for observations with low confidence levels, directly calculating cross-attention may have negative guidance on the background. To address this issue, we propose the gated cross-attention module, which is also guided by the confidence matrix. Specifically, when the observations are used as the condition of cross-attention, the background is aligned with them at first, ensuring that observations influence the background in proportion to the confidence levels. After calculating the cross-attention, the remaining proportion of the background that does not participate in the operation will be added back. Conversely, the background fully influences observations since all values in the background have relative high confidence levels generally, which can correct the observations with low confidence levels. Therefore, the confidence matrix needs to be updated accordingly, which is also implemented through a convolutional layer with $Sigmoid(\cdot)$ activation (Fig. 2c). The process of gated cross-attention can be expressed as
$x^{b}_{o}=Attention(x^{b}_{i}\odot m_{i},y_{i},y_{i})\oplus\big{(}x^{b}_{i}\odot(1-m_{i})\big{)}$ (7)
$y_{o}=Attention(y_{i},x^{b}_{i},x^{b}_{i})$ (8)
where the inputs of $Attention(\cdot)$ correspond to query ($Q$), key ($K$) and value ($V$), respectively, and $\oplus$ represents element-wise addition. The gated cross-attention module utilizes information for interactions selectively based on confidence levels, which can effectively avoid the negative impact of low-quality data.
Training loss and optimization
The mean absolute error (MAE) loss, also known as L1 loss, is employed to supervise the training of the neural network. In multivariate optimization, an issue of imbalanced optimization arises when there are significant differences in the magnitudes of losses for different variables. Therefore, it is necessary to weigh the losses for different variables. Instead of setting the weights manually, the losses for each variable (i.e., each channel) are automatically weighted to have the same magnitude in our training. The models in simulation experiments are trained for 50 epochs using the AdamW optimizer [48] and OneCycleLR scheduler [49]. The learning rate starts from 1e-6, warms to a maximum value of 1e-4 for ten epochs, and then decays gradually to 1e-10. To ensure adequate training with available GDAS observational data, the models in GDAS experiments are fine-tuned for 50 epochs based on the transfer learning strategy.
Inference speed
Compared to traditional methods, the reduction of computational consumption and execution time is a significant advantage of neural networks. In our system, Adas requires only about 0.6 seconds for the inference of data assimilation on an NVIDIA Tesla-A100 GPU, and FengWu needs about 1.5 seconds for inferring a single-step forecast. That is to say, our system can complete a periodic iteration in seconds once the observational data are available, which achieves significant efficiency improvement (about 1000-10000 times faster) compared to traditional data assimilation algorithms [50].
[…]
Acknowledgements
This research is supported by National Natural Science Foundation of China (No. 62071127, and 62101137), National Key Research and Development Program of China (No. 2022ZD0160100), Shanghai Natural Science Foundation (No. 23ZR1402900) and Shanghai Municipal Foundation (No.2021SHZDZX0103). We acknowledge the ECMWF and NCEI for their great efforts in storing and providing the invaluable data, which is very important for this work and the research community. We would also like to express our appreciation to Mr. Huihang Sun and Prof. Jiayuan Fan for their suggestion and valuable discussions during the conduction of this research, and the research team and service team in the Shanghai Artificial Intelligence Laboratory for the provision of computational resources and infrastructure.
Declarations
[…]
Competing interests The authors declare no competing interests.
[…]
Data availability The ERA5 dataset can be downloaded from the official website of Climate Data Store (CDS) at https://cds.climate.copernicus.eu/. The GDAS observational bufr files can be obtained from the Archive Information Request System (AIRS) of NCEI at https://www.ncei.noaa.gov/has/HAS.DsSelect.
[…]
Code availability The neural network model is developed using standard libraries in open-source platforms including PyTorch. Codes and model checkpoints used in this study will be available at https://github.com/kunc3301/FengWu-Adas if ready.
[…]
Author contribution L.B. and T.C. are co-corresponding authors. L.B., T.C., J.-J.L. and W.O. supervised this research. Kun Chen, L.B., F.L. and P.Y. conceived and initiated this project. Kun Chen and L.B. designed the details of model and developed the methodologies. Kun Chen and Kang Chen developed the Python code. Kun Chen, Y.X. and T.H. prepared and processed the data. Kun Chen performed experiments. Kun Chen, L.B., F.L., P.Y. and H.C. wrote the manuscript. All authors contributed to analysis of data, discussions about the results, and improvement of the presentation.
    Appendix A Extended Data
    Figure 7: RMSE and Bias skill of FengWu-Adas with random masks at each time step in simulation experiments. Rows 1 and 3 show the RMSE of the analyses for main variables over the year, and rows 2 and 4 show the Bias for them. There are no obvious performance differences compared to Fig. 3, which prove the assimilation ability and robustness of Adas. See the caption of Fig. 3 for all other details.
    Figure 8: Average RMSE skill of FengWu-Adas for all variables in simulation experiments. The different colors represent the RMSE of analyses with different observation ratios, as marked in the legend. The first subplot includes four surface variables, and other subplots correspond to five upper-air variables, as indicated in the subplot titles. The x-axis represents the single variable (and pressure level). The y-axis represents the RMSE.
    Figure 9: RMSE skill of FengWu-Adas with random initialization in simulation experiments. The different colors represent the results with different observation ratios, as marked in the bottom. Each subplot represents a single variable (and pressure level), as indicated in the subplot titles, and the subgraph provides a clearer display of the first ten days. The x-axis represents the operation time at 6-hour steps over the year. The y-axis represents the RMSE, which uses a coordinate axis with non equidistant logarithmic form to ensure visual perception. The results further demonstrate the ability of Adas to propagate sparse observation information and the stability of the system.
    Figure 10: RMSE and ACC skill of FengWu-Adas for weather forecasting over ten days in simulation experiments. The different colors represent the forecasts whose initial states are generated after assimilating observations with different ratios, as marked in the bottom. Based on the analysis produced by Adas, FengWu can still generate accurate and consistent global medium-range weather forecasts. See the caption of Fig. 6 for all other details.
    Figure 11: Average RMSE skill of FengWu-Adas for all variables in GDAS experiments. The different colors represent the RMSE of analyses after assimilating all data and non-satellite data, as marked in the legend. See the caption of Extended Data Fig. 8 for all other details.

### v4

@@
-
+Towards an end-to-end artificial intelligence driven global weather forecasting system
 Abstract
@@
 However, existing AI-based weather forecasting models rely on analysis or reanalysis products from traditional numerical weather prediction (NWP) systems as initial conditions for making predictions.
-Initial states are typically generated by traditional data assimilation components, which are computational expensive and time-consuming.
-Here we present an AI-based data assimilation model, i.e., Adas, for global weather variables.
-By introducing the confidence matrix, Adas employs gated convolution to handle sparse observations and gated cross-attention for capturing the interactions between the background and observations.
-Further, we combine Adas with the advanced AI-based forecasting model (i.e., FengWu) to construct the first end-to-end AI-based global weather forecasting system: FengWu-Adas.
-We demonstrate that Adas can assimilate global observations to produce high-quality analysis, enabling the system operate stably for long term.
-Moreover, we are the first to apply the methods to real-world scenarios, which is more challenging and has considerable practical application potential.
-We have also achieved the forecasts based on the analyses generated by AI with a skillful forecast lead time exceeding that of the IFS for the first time.
+The initial states are typically generated by traditional data assimilation components, which are computationally expensive and time-consuming.
+Here, by cyclic training to model the steady-state background error covariance and introducing the confidence matrix to characterize the quality of observations, we present an AI-based data assimilation model, i.e., Adas, for global weather variables.
+Further, we combine Adas with the advanced AI-based forecasting model (i.e., FengWu) to construct an end-to-end AI-based global weather forecasting system: FengWu-Adas.
+We demonstrate that Adas can assimilate global conventional observations to produce high-quality analysis, enabling the system to operate stably for long term.
+Moreover, the system can generate accurate end-to-end weather forecasts with comparable skill to those of the IFS, demonstrating the promising potential of data-driven approaches.
+1 Introduction
 Driven by the advancements and maturity of AI, particularly deep learning techniques, scientific intelligence has been rapidly evolving with the aim of leveraging AI to promote scientific research and discovery.
-Within the field of atmospheric science, AI has achieved remarkable achievements in various areas, such as post-processing and bias correction [1, 2, 3, 4], downscaling [5, 6], precipitation nowcasting [7, 8, 9], climate forecasting [10], and medium-range weather forecasting [11, 12, 13, 14, 15].
+Within the field of atmospheric science, AI has achieved remarkable achievements in various areas, such as post-processing and bias correction (yang2023improving; mouatadid2023adaptive; agrawal2023machine; singh2022short), downscaling (harder2022generating; chen2022rainnet), precipitation nowcasting (ravuri2021skilful; gao2022earthformer; zhang2023skilful), climate forecasting (ling2022multi), and medium-range weather forecasting (pathak2022fourcastnet; bi2023accurate; lam2023learning; chen2023fengwu; chen2023fuxi).
 Some AI-based models have demonstrated highly competitive forecasts compared to the deterministic forecasts of the state-of-the-art NWP system, i.e., the Integrated Forecasting System (IFS) from the European Centre for Medium-Range Weather Forecasts (ECMWF).
-These models are usually trained on reanalysis dataset (e.g., ERA5 [16] from ECMWF) and allow much lower computational costs and easier deployment for operational forecasting.
-Despite drawbacks such as forecast smoothness and bias drift [17], AI approaches have shown the immense potential of data-driven modeling in weather prediction, offering a new paradigm for meteorological forecasting.
-However, the AI-based weather forecasting models still require analysis products generated through the process of data assimilation in the traditional NWP system for making predictions (Fig.
-1).
-Specifically, data assimilation aims to obtain the best estimate of the true state of the Earth system (known as the analysis) and provide an accurate initial state for weather prediction, thereby improving the forecast performance.
-With periodic corrections applied to the forecasts, data assimilation allows for a more stable and accurate long-term prediction, which is a critical component in global weather forecasting system.
-This process relies heavily on observations as they provide crucial information that closely represents the true state of the atmosphere.
-For example, the earliest initial conditions were obtained by interpolating observations onto the grid points of the grid space [18].
-Modern data assimilation techniques are usually achieved by integrating observations with the short-range forecasts (i.e., the background), primarily including two categories: Kalman filters and variational methods [19].
+These models are usually trained on reanalysis dataset (e.g., ERA5 (hersbach2020era5) from ECMWF) and allow much lower computational costs and easier deployment for operational forecasting.
+Despite drawbacks such as forecast smoothness, bias drift and lack of physical consistency (ben2023rise; bonavita2024some), AI approaches have shown the immense potential of data-driven modeling in weather prediction, offering a new paradigm for meteorological forecasting.
+However, the AI-based weather forecasting models still require analysis products generated through the process of data assimilation in the traditional NWP system for making predictions (Figure 1).
+Specifically, data assimilation aims to obtain the best estimate of the true state of the Earth system (known as the analysis) and provide an accurate initial state for weather prediction, thus improving the forecast performance.
+In the operational weather forecasting system, data assimilation is a critical component that ensures the long-term stable operation of the system by periodically correcting the forecast.
+Beyond the predictability limit, which is believed to be about 15 days, the forecast results without data assimilation will become completely unreliable.
+This process is heavily dependent on observations, as they provide crucial information that closely represents the true state of the atmosphere.
+For example, the earliest initial conditions were obtained by interpolating observations onto the grid points of the grid space (richardson1922weather).
+Modern data assimilation techniques are usually achieved by integrating observations with short-range forecasts (i.e., the background), primarily including two categories: Kalman filters and variational methods (rabier2003variational).
 The main motivation for introducing the background, or the first guess, lies in the incomplete nature of observations.
-The sparsity of observations makes it difficult to obtain a global estimate, resulting in low-quality interpolation analysis.
-Besides, not all state variables can be observed directly.
-Currently, AI methods have primarily been applied to specific processes within data assimilation to address the limitations of traditional data assimilation techniques [20, 21].
-For instance, implicit neural representations (INRs) [22] and various autoencoders (AEs) [23, 24, 25] have been employed as reduced-order-models (ROMs) to tackle high-dimensional challenges.
-These methods replace classical linear models such as Proper Orthogonal Decomposition (POD) [26, 27] in latent assimilation, providing an efficient framework for the representation and reconstruction of state variables in the latent space.
-Neural networks have also been utilized to derive tangent-linear and adjoint models in 4D variational (4D-Var) data assimilation [28], provide localization functions for ensemble Kalman filters (EnKF) [29] and estimate error covariance matrix [30, 31].
-Moreover, there exist strong mathematical similarities between machine learning and data assimilation, particularly the variational data assimilation [32], which enables the optimization of the cost function in variational data assimilation using auto-differentiation [24, 33].
-However, these studies have not involved altering the algorithms of data assimilation fundamentally.
-While traditional data assimilation algorithms are underpinned by rigorous mathematical theory and physical priors, they inevitably rely on certain assumptions as prerequisites.
-In practical operational applications, they often face constraints due to high computational costs, necessitating various approximations for solution [34].
-In situations where supervised information is available, neural networks are expected to capture the correlations among data during the assimilation automatically and provide a new kind of algorithms [35].
-Efforts have been made in some simple scenarios [36, 37, 38], but they simply feed the data into networks without any specialized design.
-And so far, there have been no attempts in data assimilation for real observational data in multivariate global weather scenarios.
-In this study, we present Adas, a novel data assimilation model for global weather variables, which can assimilate sparse observations with different qualities and provide initial conditions for weather forecasting.
-Inspired by the gating mask in gated convolution [39] and the error covariance matrix in traditional data assimilation techniques, we introduce the confidence matrix to characterize the availability and quality of observations during the data assimilation.
+The sparsity and irregularity of observations make it difficult to obtain a global estimate, resulting in low-quality interpolation analysis.
+Besides, not all instruments can observe geophysical variables directly.
+Figure 1: The progression of global weather forecasting system.
+Traditional NWP systems consist of physical weather forecasting model and data assimilation.
+The breakthrough of AI-based medium-range weather forecasting models occurred in 2022-2023 with highly competitive performance in terms of accuracy, but they still rely on the NWP systems for making predictions.
+Our work is dedicated to exploring the possibility of an end-to-end global weather forecasting system which is driven purely by AI.
+Currently, AI methods have been applied primarily to specific processes within data assimilation to address the limitations of traditional data assimilation techniques (buizza2022data; cheng2023machine).
+For example, implicit neural representations (INRs) (li2023latent) and various autoencoders (AEs) (peyron2021latent; melinc2023neural; amendola2021data) have been employed as reduced-order models (ROMs) to address high-dimensional challenges.
+These methods replace classical linear models such as proper orthogonal decomposition (POD) (cheng2023generalised; pawar2022equation) in latent assimilation, providing an efficient framework for the representation and reconstruction of geophysical variables in latent space.
+Neural networks have also been utilized to derive tangent-linear and adjoint models in 4D variational (4DVar) data assimilation (hatfield2021building), provide localization functions for ensemble Kalman filters (EnKF) (wang2023convolutional) and estimate error covariance matrix (cheng2022observation; penny2022integrating).
+Moreover, there exist strong mathematical similarities between machine learning and data assimilation, particularly the variational data assimilation (geer2021learning), which enables the optimization of the cost function in variational data assimilation using auto-differentiation (melinc2023neural; xiao2023fengwu).
+However, these studies have not involved fundamentally altering the algorithms of data assimilation.
+While traditional data assimilation algorithms are underpinned by rigorous mathematical theory and physical models, they inevitably rely on certain assumptions as prerequisites.
+In practical operational applications, they often face constraints due to high computational costs, which requires various approximations for solution (carrassi2018data).
+In situations where supervised information is available, neural networks are expected to automatically capture the correlations among data during the assimilation and provide a new alternative algorithm (fablet2021learning).
+Efforts have been made in idealized data assimilation experiments (rozet2023score; huang2024diffda; chen2024fnp), and further research focuses on data-driven end-to-end weather forecasting systems, broadly on systems that can initialize forecasts without needing traditional physically-based analysis (vaughan2024aardvark; xu2024fuxi; mcnally2024data).
+Some follow the paradigm of traditional NWP systems and specially design data assimilation models to work together with AI-based weather forecasting models in a forecast-assimilation cycle to form a system (xu2024fuxi).
+Others implicitly integrate the process of data assimilation and weather forecasting, using a single model to make predictions directly from observations (vaughan2024aardvark; mcnally2024data).
+In this study, we present Adas, a novel data assimilation model for global weather variables, which can assimilate sparse conventional observations with different qualities and provide initial conditions for weather forecasting.
+The key of data assimilation lies in adjusting the weights of the background and observations based on error covariance, and traditional data assimilation techniques primarily derive optimal analysis under the Bayesian framework.
+In our method, we model the steady-state (i.e., optimal dynamic balance between forecasting error growth and assimilation error reduction) background error covariance by cyclic training and introduce the confidence matrix to characterize the availability and quality of observations during the data assimilation.
 With the guidance of confidence matrix, Adas employs the gated convolution module to handle sparse observations and the gated cross-attention module for capturing the interactions between the background and observations efficiently.
-The prediction of the advanced AI-based weather forecasting model, FengWu [14], is used to generate the background for data assimilation.
-In return, the analysis of Adas is then used as the initial state of FengWu for making predictions at the next time step, thus forming the first end-to-end AI-based global weather forecasting system: FengWu-Adas.
-Fig.
-1 shows the progression of global weather forecasting system and the main motivation of this research.
+The prediction of the advanced AI-based weather forecasting model, FengWu (chen2023fengwu), is used to generate the background for data assimilation.
+In return, the analysis of Adas is then used as the initial state of FengWu for making predictions at the next time step, thus forming an end-to-end AI-based global weather forecasting system: FengWu-Adas.
 We demonstrate that our system can operate stably for long term, and our method is robust and flexible to be combined with any state-of-the-art weather forecasting model without retraining or fine-tuning.
-The simulation experiments have proven the superior performance of FengWu-Adas in capturing the distribution patterns of the background error and producing high-quality analysis.
-And we further apply FengWu-Adas to real-world scenarios, which is more challenging and has considerable practical application potential.
-We have also achieved the forecasts based on the analyses generated by AI with a skillful forecast lead time of more than 8.75 days, which exceeds that of the IFS for the first time.
+The ideal experiments have demonstrated the superior properties of FengWu-Adas in producing accurate analysis.
+Further, we apply FengWu-Adas to real-world scenarios, conducting a comprehensive and detailed evaluation to the system and making comparison with the IFS.
+The results show that FengWu-Adas can generate high-quality analysis and accurate end-to-end global weather forecasts although it only uses conventional observations, demonstrating the promising potential of data-driven approaches.
 Results
 Principle of FengWu-Adas
-The framework of FengWu-Adas is shown schematically in Fig.
-2a.
+The framework of FengWu-Adas is shown schematically in Figure 2a.
 For the given initial state denoted as $\dot{x}_{t}$ at time $t$, FengWu can make multi-step predictions $\hat{x}_{t+\Delta t}$, $\hat{x}_{t+2\Delta t}$, …, $\hat{x}_{t+k\Delta t}$ at time $t+\Delta t$, $t+2\Delta t$, …, $t+k\Delta t$ in an auto-regressive manner, where $k$ denotes the time index and $\Delta t=6h$ is the temporal spacing of single-step prediction.
 Considering that the accuracy of the background will directly affect the quality of the analysis, the single-step prediction of FengWu is used as the background $x_{t+\Delta t}^{b}$ for data assimilation.
-After obtaining the observations $y_{t+\Delta t}$ at the corresponding time, confidence levels ranging from 0 to 1 can be assigned to each datum based on the availability and quality of the observation, thereby generating the confidence matrix.
-The confidence matrix $m$, serving as supplementary information, is sent into Adas along with observations for assimilation.
+After obtaining the observations $y_{t+\Delta t}$ at the corresponding time, confidence levels ranging from 0 to 1 can be assigned to each value in observation field based on the availability and quality of the observation, thereby generating the confidence matrix.
+The confidence matrix $m$, serving as additional information for the uncertainty of observation error, is sent into Adas along with observations for assimilation.
+For Adas, the input dataset is the background and observations, while the target dataset is ERA5, and the learning objective is to integrate the background forecast and observations to obtain analysis close to ERA5.
 Subsequently, the analysis $x_{t+\Delta t}^{a}$ produced by Adas will be the initial state for the next prediction, i.e., the input of FengWu at next time and so on, in a cyclic manner.
-The multi-step forecasts are predicted through auto-regression based on the analysis.
-Adas utilizes a dual encoder composed of gated interaction blocks to extract features of the background and sparse observations separately and capture the interactions between them (Fig.
-2b).
+Different from the general single-step training with a fixed lead background, this cyclic training makes the model more robust to the background, because the background quality by starting the forecast with analysis at the previous time changes dynamically with the convergence of the model training.
+By continuously iterating to optimize the implicit modeling of background error covariance, the system can achieve a steady state that balances the error growth of forecasting and the error reduction of assimilation.
+This approach is not limited to the optimal analysis of single-step assimilation, instead seeking a dynamic equilibrium for the forecasting and assimilation of the entire system, whose superiority has been previously substantiated (yang2006data; kalnay20074).
+As an end-to-end global weather forecasting system, FengWu-Adas can also perform multi-step forecasts through auto-regression based on the analysis.
+Figure 2: The framework of FengWu-Adas and network architecture of Adas.
+a, The framework and process of FengWu-Adas.
+The prediction of FengWu provides the background for data assimilation and the analysis of Adas serves as the initial state for weather forecasting.
+The parameters of FengWu are frozen during training and not further fine-tuned or involved in joint optimization.
+b, Overview of Adas’s network structure.
+The dual encoder composed of gated interaction blocks extract multi-scale features of the background and sparse observations separately and capture the interactions between them, and then the features are fused and decoded.
+c, Details of dual gated interaction block.
+The gated convolution and gated cross-attention modules are designed for sparse observations, which are both guided by the confidence matrix.
+d, Simple schematic diagram of patch merging and patch expanding.
+The patch merging performs the rearrangement operation and linear layer to achieve down-sampling and the patch expanding performs the opposite operation to achieve up-sampling.
+Adas utilizes a dual encoder composed of gated interaction blocks to extract features of the background and sparse observations separately and capture the interactions between them (Figure 2b).
 All the inputs with a shape of $H_{0}\times W_{0}\times C_{0}$ are converted into a unified latent shape of $D\times H\times W\times C$ through the patch embedding module so that the network can capture their spatial relationships in the 3D mesh.
@@
 $H$ and $W$ correspond to the number of grid points in the longitudinal and latitudinal directions after embedding, and $D$ represents the number of levels in the vertical direction.
-By performing down-sampling and up-sampling through the patch merging and patch expanding module (Fig.
-2d), the encoder can capture multi-scale meteorological features.
-The dual gated interaction block extracts features from the background and observations with standard convolution and gated convolution, along with efficient information interactions through the gated cross-attention module (Fig.
-2c).
+By performing down-sampling and up-sampling through the patch merging and patch expanding module (Figure 2d), the encoder can capture multi-scale meteorological features.
+The dual gated interaction block extracts features from the background and observations with standard convolution and gated convolution, along with efficient information interactions through the gated cross-attention module (Figure 2c).
 Both the gated convolution and gated cross-attention are guided by the confidence matrix, which is updated together with the features to represent the availability and quality of observations (details in Methods).
-After the feature fusion block implemented by a convolutional layer, the features are sent into a series of Transformer [40, 41] blocks and then recovered to the original size.
+After the feature fusion block implemented by a convolutional layer, the features are sent into a series of Transformer (vaswani2017attention; dosovitskiy2020image) blocks and then recovered to the original size.
 The convolutional operation can bring inductive bias of locality, and the attention mechanism enables the model to capture long-range dependencies.
-The attention block utilizes the shifted-window mechaism [42] adjusted according to geographical rules to reduce the computational cost and ensure continuity in the latitudinal direction, and $N=8$ in our implementation.
-Superior performance of FengWu-Adas in simulation experiments
-For testing the performance of FengWu-Adas, the Observing System Simulation Experiments (OSSEs) are performed by utilizing the ERA5 dataset firstly.
-The experiment proceeds with alternating weather forecasting and data assimilation steps: the single-step forecast provides the background for data assimilation, and the analysis is used as the initial state for the next forecasting step.
-ERA5 is regarded as the ground truth, and the simulated observations are obtained by adding a mask to the ground truth.
-The mask is used as the simulated confidence matrix for data assimilation.
-Therefore, the confidence matrix in simulation experiments is binary, with 1 indicating that the observation in corresponding grid point is available and accurate, and 0 indicating unavailable observation.
-The confidence matrix in real scenarios will be more complex and the confidence value is produced based on the quality of the observations.
-To improve the robustness of observation locations, the simulated masks for training are generated randomly with a fixed mask ratio at each time step.
-However, most observation locations are generally fixed in real-world scenarios, and the same strategy for testing not only deviates from reality but also poses potential issues in the cyclic experiments.
-Specifically, a random mask implies that any position has the probability of being observed.
-When the number of steps in the cyclic experiments is sufficiently large, the observation positions would approximately cover the whole world, which is clearly unrealistic.
-To address this concern, the masks for testing are fixed even though we find that this has almost no impact on performance in our experiments (see the comparison between Fig.
-3 and Extended Data Fig.
-7).
-Here, we conduct three sets of experiments with different observation ratios, i.e., 90%, 99% and 99.9%.
-In other words, the ratios of available observations to the total number of grid points are 10%, 1% and 0.1%.
-The system is initialized using the ERA5 data at 2018-01-01 00:00UTC, which serves as the initial state for the first-step forecasting.
-Once the system is initialized, ERA5 data is invisible throughout the whole year, except for generating simulated observations.
-Fig.
-3 illustrates the variations in RMSE and Bias (defined in Methods) of the analyses for main variables over the year.
-Both RMSE and Bias are computed against ERA5.
-The results with different observation ratios are represented by different colors, as marked in the bottom.
-To enhance clarity, the original data is presented with reduced opacity, while the solid lines indicate the smoothed values after the exponential moving average (EMA) with a smoothing factor of 90%.
-The RMSE of the analyses for all variables with all observation ratios is maintained at a low level for the whole year, showing the long-term stability of the system and the ability to generate accurate analysis based on Adas.
-The dashed lines are the RMSE of 6-hour forecast of the IFS, which represent fairly advanced error levels.
-Even when the observation ratio is only 0.1% (corresponding to an absolute number of observations of less than 0.1 million), the RMSE for all variables is lower than the dashed lines at almost all times.
-As the observation ratio grows, the RMSE of the analysis decreases gradually, and the Bias remains at around 0 more stably, revealing the system’s potential to eliminate bias drift of AI-based forecasting models and maintain conservation.
-The average RMSE over the whole year for all variables is shown in Extended Data Fig.
-8.
-The visualizations for z500 during the data assimilation, along with the error distributions and analysis increments, are presented in Fig.
-4.
-The visualization date-time is randomly selected at 2018-01-26 06:00UTC, where the background is obtained from FengWu by performing single-step forecasting of the analysis at 2018-01-26 00:00UTC.
-The first row displays the visualizations of ERA5 (i.e., the ground truth, Fig.
-4a) and the background (Fig.
-4b), and the error distributions of the background (i.e., the background minus ERA5, Fig.
-4c).
-The other rows show the visualizations of random simulated observations (the first column), the analysis increments (i.e., the analysis minus the background, the second column), and the error distributions of the analysis after data assimilation (i.e., the analysis minus ERA5, the last column) with different observation ratios.
-All the analysis increments exhibit a distribution similar to the background errors, indicating that Adas captures the error distribution of the background effectively.
-And all the errors of the analysis are reduced to varying degrees after data assimilation.
-As expected, the analysis increment is closer to the background error and the analysis error is smaller with the increase of observation information.
-In order to further demonstrate the stability of the system, a completely random initial state is used to start the system.
-Extended Data Fig.
-9 shows the corresponding RMSE, and the subgraph provides a clearer display of the first ten days.
-The y-axis in these graphs uses a coordinate axis with non equidistant logarithmic form to ensure visual perception.
-The first-step forecast after random initialization will produce a terrible performance, but all the errors converge quickly to a stable state with the same levels of metrics as that of starting the system with ERA5 in a few days.
-And the convergence speed accelerates as the number of observations increases, which is reasonable.
-This indicates that Adas can perform data assimilation on the background with any quality, ensuring that the system can correct back to the stable state in a short period of time without affecting subsequent operations even if it encounters a malfunction that leads to complete derailment of the forecast.
-As an end-to-end global weather forecasting system that can operate independently based on observational data, FengWu-Adas can generate multi-step forecasts on its own analysis (Fig.
-2a).
-The average RMSE and ACC skill for weather forecasting of FengWu-Adas over ten days is shown in Extended Data Fig.
-10.
-It should be emphasized that FengWu is trained on ERA5 data [14] and is not fine-tuned to our task.
-Therefore, our method is flexible to be combined with any state-of-the-art weather forecasting model.
-Actually, our pre-trained model can assimilate observations on arbitrary background without retraining or fine-tuning, as shown in Extended Data Fig.
-9.
-The universal capability allows Adas to be directly integrated with other weather forecasting models as a plug-and-play component, and the forecasting skill of the system will improve as the forecasting model develops.
+The attention block utilizes the shifted-window mechanism (liu2021swin) adjusted according to geographical rules to reduce the computational cost and ensure continuity in the latitudinal direction.
+Superiority of FengWu-Adas in ideal experiments
+The ideal cyclic forecast-assimilation experiments are first conducted using the ERA5 dataset to verify the performance and characteristics of FengWu-Adas.
+The experiments proceed with alternating weather forecasting and data assimilation steps: the single-step forecast serves as the background for data assimilation, and the analysis provides the initial state for the next forecasting step.
+ERA5 is considered as the ground truth, and simulated observations are obtained by sampling it with binary mask, which is then used as the confidence matrix for data assimilation.
+For simplicity, all variables are sampled independently in a certain proportion.
+Simulated masks for training are generated randomly at each time step to enhance robustness with respect to the observation locations.
+During the testing period, however, the masks remain fixed to avoid spurious performance caused by variations in information coverage at different times.
+The system is initialized with a 5-day lead forecast as the background, and except for producing simulated observations, ERA5 data is invisible throughout the entire experiments.
+Figure 3a illustrates the RMSE (defined in Methods) variations of the analyses for $z500$ and $t850$ variables over the year, and the RMSE is calculated against ERA5.
+The results with different observation ratios are represented by different colors, as marked in the legend.
+For clarity, the original data is presented with reduced opacity, while solid lines indicate the smoothed values after exponential moving average (EMA).
+The RMSE of the analyses with all observation ratios rapidly decreases and remains at a low level for the whole year, showing the long-term stability of the system and its ability to generate accurate analysis based on Adas.
+It should be emphasized that FengWu is trained on ERA5 data (chen2023fengwu) and is not fine-tuned to our task.
+As the observation ratio (i.e., the amount of observation information) increases, the speed at which the analysis RMSE decreases becomes faster, and the analysis RMSE in the final steady state is also lower, which is consistent with our general understanding.
+Figure 3: The performance and properties of FengWu-Adas in ideal experiments.
+The RMSE (lower is better) is calculated against ERA5.
+a, RMSE variations of the analyses for $z500$ and $t850$ variables with different observation ratios in a whole year.
+The forecast with a 5-day lead time is used as the initial background to start the system.
+As the system continuously executes the cyclic forecast-assimilation, the RMSE of the analysis decreases rapidly and converges to a lower level.
+b, The performance differences between cyclic training and single-step training.
+The cyclic training directly models the background error covariance at steady state, circumventing the issue of error accumulation present in single-step training during cyclic forecast-assimilation experiments and achieving lower steady-state errors.
+c, The performance when starting the system with completely random Gaussian noise as the initial background.
+The system can still quickly recover to normal steady-state levels, showing outstanding stability and generalization capability to assimilate observations on arbitrary background.
+d, The power spectra of the 10-day lead background and analysis for $z500$ variable.
+After data assimilation, Adas can significantly correct the spectral degradation problem of FengWu forecasts with long lead time at small and medium scales.
+e, The sensitivity of the system to the noises with varying magnitudes.
+The horizontal axes correspond to different random Gaussian noise proportions of 0, 0.1%, 0.5%, 1%, 2.5%, 5%, and 10%, respectively.
+It demonstrates the robustness and anti-interference capability of Adas in the presence of noise.
+We compare the RMSE variations of the analyses under single-step training and cyclic training to demonstrate the effectiveness of the training method in Figure 3b.
+The cyclic training directly models the background error covariance at steady state, circumventing the issue of error accumulation present in single-step training during cyclic forecast-assimilation experiments and achieving lower steady-state errors.
+Moreover, as the model encounters a more diverse range of background samples during training, it exhibits enhanced robustness to backgrounds with varying qualities and results in reduced error fluctuations.
+With this favorable property, even when starting the system with completely random Gaussian noise as the initial background, it can converge rapidly to the same steady-state level within a few days, and the convergence speed accelerates as the number of observations increases (Figure 3c).
+This ensures that the system can correct back to the stable state in a short period of time without affecting subsequent operations even if it encounters a malfunction that leads to complete derailment of the forecast.
+The generalization capability allows Adas to assimilate observations on arbitrary background without retraining or fine-tuning, so it is flexible to be integrated with any state-of-the-art weather forecasting model as a plug-and-play component.
+A major challenge of current AI-based forecasting models is the smoothness of their predictions, which often manifests itself as degraded spectra (kochkov2024neural; lang2024aifs).
+Preserving fine-scale information is also important for data assimilation, allowing for better representation of errors at small and medium scales, which then grow to affect larger scales.
+Figure 3d shows the power spectra of the 10-day lead background and analysis for $z500$ variable.
+After data assimilation, Adas can significantly correct the spectral degradation problem of FengWu forecasts with long lead times at small and medium scales.
+And as the amount of observation information increases, the spectra of analysis will be closer to that of ERA5.
+In addition, we examine the sensitivity of the system to the noises with varying magnitudes and the balance problem when assimilating only temperature observations.
+Figure 3e shows the changes in RMSE of $z500$ and $t850$ with random Gaussian noise proportions of 0, 0.1%, 0.5%, 1%, 2.5%, 5%, and 10%, respectively.
+The change in analysis error caused by observational error is much smaller than the observational error itself, demonstrating the robustness and anti-interference capability of Adas in the presence of noise.
+As the number of observations increases, the model’s sensitivity to observational error will decrease further.
 Applying FengWu-Adas to real-world scenarios
-The data assimilation for real observational data is more challenging than simulation experiments and has considerable practical application potential.
-To validate the performance in real-world scenarios, we apply FengWu-Adas to the assimilation for observations of Global Data Assimilation System (GDAS).
-The values of state variables are interpolated onto the grid points to generate the observations, and the confidence matrix is produced according to the availability and quality of the observations.
-Observing that satellite-derived data often have larger systematic errors, we conducted two sets of experiments to assimilate all data and non-satellite data of GDAS.
-Fig.
-5 illustrates the variations in RMSE and Bias of the analyses for main variables over the year, where the black crosses indicate that the observational data at the corresponding time are missing.
-When FengWu-Adas is applied to the assimilation for real observational data, the system can still operate stably for long term, and the RMSE of some variables such as specific humidity can reach a level comparable to that of simulation experiments and the IFS (Extended Data Fig.
-11).
-The results show that although satellite-derived data are less accurate, they can provide more information to improve the quality of the analysis, and it is expected that neural networks are robust to noise.
-This holds true for most variables except for the pressure level of 50hPa, which may be caused by the small number of high-altitude observational data.
-What’s more, Adas is robust to handle the lack of observations at a certain time.
-When the input observations is an all-zeros matrix, it can perform a function like post-processing to maintain the operation and produce the analysis and forecasts.
-The average RMSE skill for weather forecasting based on the analyses over ten days is shown in Fig.
-6, and the intersections with dashed line correspond to the skillful forecast lead times with ACC=0.6 for z500.
-The analyses generated after assimilating all data and non-satellite data are used as the initial states for weather forecasting, and the skillful forecast lead times reach 8.75 days and 8.25 days, respectively.
-The results reflect the forecasting skill of our system after assimilating real observational data, and we have achieved the forecasts based on the analyses generated by AI with a skillful forecast lead time exceeding that of the IFS (8.5 days [43]) for the first time.
-The GDAS experiments demonstrate the stability and considerable potential for practical applications of our system.
+Data assimilation for real observational data is more challenging than simulation experiments and has considerable practical application potential.
+To validate the performance in real-world scenarios, we apply FengWu-Adas to the assimilation for conventional observations from the Global Data Assimilation System (GDAS).
+The main data types and information of GDAS observations we use are given in Table 1.
+The values of geophysical variables are processed into grid data through nearest neighbor interpolation to generate the observation fields, and the confidence matrix is produced according to the availability and quality of the observations (details in Methods).
+Figure 4a and Figure 4b illustrate the RMSE variations of the analyses for $z500$ and $t850$ variables over the year in the cyclic forecast-assimilation experiment, which are calculated against ERA5 and the 200 reserved GDAS observation columns, respectively.
+When FengWu-Adas is applied to the assimilation for real observational data, the system can still operate stably for long term.
+Compared with the baseline of 3D variational (3DVar) data assimilation, Adas has lower analysis error and stronger stability.
+It is worth noting that the 3DVar we reproduced may not be optimal, and it may achieve improved performance with careful tuning.
+It is encouraging that the analysis quality of Adas evaluated at the stations shows a comparable level to IFS-Analysis and ERA5, although Adas only assimilates conventional observations.
+Furthermore, we analyze the statistical characteristics of the RMSE across the evaluation stations at all times and draw the box plot in Figure 4c.
+Whether considering the median, interquartile range (IQR) or outliers, Adas shows reasonable results and is not significantly worse than IFS-Analysis and ERA5.
+Figure 4d shows the scatter plot and corresponding fits between the observation and analysis values across the evaluation stations at all times, with the coefficient of determination $R^{2}$ and the slope of the fitted line.
+The analysis values of Adas fit the observations well, and its $R^{2}$ is only slightly lower than that of IFS-Analysis and ERA5.
+Figure 5a and Figure 5b visualize the fields of $z500$ and $t850$ variables and their error distributions with ERA5 during data assimilation, respectively, where the background uses the $96h$ lead forecast based on ERA5 and the date-time is randomly selected at 2017-01-27 00:00 UTC.
+Based on the information provided by observations, Adas can effectively reduce the background error, and the analysis in areas with dense observations will be more accurate.
+Figure 5c visualizes the RMSE distribution evaluated on the reserved stations at the same date-time, they all show similar errors since the learning target of Adas is ERA5.
+Table 1: Main data information of GDAS observations used in FengWu-Adas system.
+GDAS observational data consists of many messages with different types, each of which contains many observation columns with different height levels.
+The conventional types of observations in GDAS are used for data assimilation and evaluation in our experiments, covering the main geophysical variables such as temperature, humidity and wind speed.
+The table shows the number of messages and observation columns of each data type at 2017-01-01 00:00 UTC.
+Upper-air observations are mainly provided by ADPUPA and AIRCFT, and each observation column of ADPUPA usually contains 1 to more than 100 unfixed height levels.
+Surface observations of land and ocean are mainly provided ADPSFC and SFCSHP, respectively.
+Type Number of Number of Description
+Messages Obs. Columns
+ADPUPA 338 1403 Upper-Air (RAOB, PIBAL, RECCO,
+DROPS) Reports
+AIRCFT 21 2255 AIREP/PIREP, AMDAR(ASDAR/ACARS),
+E-ADAS(AMDAR BUFR) ACFT
+ADPSFC 1117 122093 Surface Land (SYNOP, METAR) Reports
+SFCSHP 150 20268 Surface Marine (SHIP, BUOY,
+C-MAN Platform) Reports
+Figure 4: The performance of FengWu-Adas to produce analysis with GDAS real observations.
+a, RMSE variations of the analyses for $z500$ and $t850$ variables evaluated on ERA5 and the comparison with 3DVar algorithm.
+By alternating between forecasting and data assimilation, the system can also maintain long-term stability in real-world scenarios.
+b, RMSE variations of the analyses evaluated on 200 reserved GDAS observation columns and comparison with IFS-Analysis and ERA5.
+The analysis quality of Adas evaluated at the stations shows a comparable level to IFS-Analysis and ERA5, although Adas only assimilates conventional observations.
+c, The box plot of the RMSE across the evaluation stations at all times.
+Whether considering the median, IQR or outliers, Adas shows reasonable results and is not significantly worse than IFS-Analysis and ERA5.
+d, The scatter plot and corresponding fits between the observation and analysis values across the evaluation stations at all times.
+Figure 5: Visualization of the fields during data assimilation and the RMSE distribution evaluated on the reserved stations.
+a, Visualization of $z500$ variable and corresponding error distributions with ERA5 during data assimilation.
+The date-time is randomly selected at 2017-01-27 00:00 UTC and the background uses the $96h$ forecast based on ERA5.
+b, Visualization of $t850$ variable and error distributions during data assimilation.
+c, Visualization of the RMSE distribution at the reserved stations and comparison with IFS-Analysis and ERA5.
+The date-time is also selected at 2017-01-27 00:00 UTC, and the analysis of Adas shows similar errors to IFS-Analysis and ERA5.
+To comprehensively evaluate the end-to-end forecasting skill of FengWu-Adas, we test its performance on ERA5, GDAS, and the Integrated Global Radiosonde Archive (IGRA) datasets and compare it with the IFS-HRES.
+IGRA covers radiosonde observations from more than 2,800 stations around the world (vandal2024global).
+The evaluation on GDAS is calculated against 1000 stations with highest quality to avoid the impact of poor observations.
+Figure 6a shows the annual average RMSE skill for $z500$ and $t850$ variables over 10 days.
+All evaluations show the same trend, with FengWu-Adas having higher forecast RMSE than IFS-HRES in the early stage and lower than IFS-HRES starting around 7-8 days.
+Since Adas assimilates far fewer observations than the IFS, resulting in less accurate analysis, FengWu-Adas has a larger RMSE for forecasts at shorter lead times.
+However, FengWu accumulates errors more slowly for multi-step forecasts, so FengWu-Adas has a smaller RMSE for forecasts at longer lead times.
+Adas mainly provides the initial state for the forecasting model, which contributes more at shorter lead times, and FengWu contributes more at longer lead times.
+Figure 6b shows the box plot of the prediction RMSE with different lead times.
+Similar to the results of average RMSE in Figure 6a, FengWu-Adas performs slightly worse than IFS-HRES at shorter lead times, but performs better at longer lead times, with lower medians and IQRs.
+The scatter plot and corresponding fits between the observation and prediction values across the evaluation stations are illustrated in Figure 7.
+The first two rows and last two rows correspond to the $z500$ and $t850$ variables, respectively.
+The columns have different lead times of 1, 5 and 10 days, and the coefficient of determination $R^{2}$ and the slope of the fitted line are marked in the subplot.
+The prediction of FengWu-Adas is more consistent with the observations at longer lead times and has a higher $R^{2}$.
+In addition, we visualize the distribution of forecast RMSE at the GDAS evaluation stations in Figure 8.
+The initialization date-time is also selected as 2017-01-27 00:00 UTC, which is consistent with Figure 5.
+The first, second and third columns correspond to the RMSE distributions with 1-day, 5-day and 10-day forecast lead times, respectively.
+It can be seen that in the forecast results with 10-day lead time, IFS-HRES has a large deviation in North America, while FengWu-Adas has a better forecast performance in this region.
+In general, FengWu-Adas shows accurate end-to-end global weather forecasts, demonstrating the promising potential of data-driven approaches.
+Figure 6: The performance of FengWu-Adas to produce end-to-end global weather forecasts and the comparison with IFS-HRES.
+a, Annual average RMSE skill for $z500$ and $t850$ variables over 10 days evaluated on ERA5, GDAS and IGRA datasets.
+All evaluations show the same trend, with FengWu-Adas having higher prediction RMSE than IFS-HRES in the early stage and lower than IFS-HRES starting around 7-8 days.
+FengWu-Adas shows accurate end-to-end global weather forecasts, demonstrating the promising potential of data-driven approaches.
+b, The box plot of the prediction RMSE across the GDAS evaluation stations with different lead times.
+Similar to the results of average RMSE, FengWu-Adas performs slightly worse than IFS-HRES at shorter lead times, but performs better at longer lead times with lower medians and IQRs.
+Figure 7: The scatter plot and corresponding fits between the observation and prediction values across the evaluation stations.
+The first two rows and last two rows correspond to the $z500$ and $t850$ variables, respectively.
+The columns have different lead times of 1, 5 and 10 days, and the coefficient of determination $R^{2}$ and the slope of the fitted line are marked in the subplot.
+The prediction of FengWu-Adas is more consistent with the observations at longer lead times and has a higher $R^{2}$.
+Figure 8: Visualization of the RMSE comparison of FengWu-Adas and IFS-HRES forecasts for $z500$ and $t850$ variables at the GDAS stations.
+The initialization date-time is set at 2017-01-27 00:00 UTC, which is consistent with Figure 5, and the RMSE is calculated against GDAS observations.
+The columns have different lead times of 1, 5, and 10 days, and the first two rows and last two rows correspond to the $z500$ and $t850$ variables, respectively.
+In the forecast results with 10-day lead time, IFS-HRES has a large deviation in North America, while FengWu-Adas has a better forecast performance in this region.
 Discussion
-There are two main reasons that make assimilating real observational data more challenging than simulated observations.
-On one hand, the distribution patterns of observational data are different between these scenarios.
-In simulation experiments, observations generated through random sampling or other methods are approximately distributed uniformly, which is an ideal distribution pattern that yields favorable performance.
-However, observational data often exhibit a severely non-uniform distribution due to various geographical and human factors in real-world scenarios.
-For instance, there is a scarcity of observations in polar regions in most cases, which can significantly impact the regional analysis.
-Such impacts may be propagated through the system’s operation, potentially leading to error accumulation or even collapse of the system.
-On the other hand, there exists an inherent inconsistency between observational values in the observational space and grid-point values in the grid space.
-Observational values represent the true and objective reflection of individual states at specific points in space, while grid-point values represent the overall states of a region.
-Therefore, observational values tend to exhibit larger fluctuations, while grid-point values are more stable and smooth theoretically.
-This inconsistency can be alleviated as the resolution of grid space increases and the quantity of observational data grows, but it remains a problem that cannot be ignored at present.
-Traditional data assimilation algorithms address this inconsistency by aligning the grid space to the observational space using observation operator.
-In contrast, our method requires an operator for the inverse mapping, which may be achieved by adding a pre-processing module instead of simple interpolation and further enhance the performance of our method.
-In summary, there exists a natural gap between simulated experiments and real-world systems.
-While simulation experiments are necessary for academic research and algorithm improvement, the ability to be extended to real-world scenarios is of paramount importance for practical applications.
-We have achieved stable long-term assimilation of real observational data for the first time, which can help forecasting models get rid of the dilemma of relying on the initial states provided by inefficient traditional data assimilation components, and promote the rapid development of research in related fields.
-The future work will mainly focus on a more universal and high-performance framework.
-For example, techniques based on continuous space modeling [44] may provide a train of thought for better solution, which can work directly with the off-grid real observations without pre-processing or observation operator that may introduce additional errors.
-Nowadays, high-quality data has always been essential for the development of AI and many other technologies, and the reanalysis datasets still hold an irreplaceable significance.
-Therefore, developing physical models and traditional algorithms remains crucial.
-The purpose of our study is not to replace the traditional NWP systems but to demonstrate AI methods’ potential in tackling real-world challenges.
-Both physical techniques and AI methods have their own strengths and limitations, and they should be treated equally and considered as powerful tools for benefiting science and society.
+Significant progress has been made in AI-based weather forecasting research on the ideal setting based on ERA5 dataset, which prompts our focus to shift towards end-to-end forecasting based on realistic observations.
+Some ambitious yet radical endeavors have aimed to discard the background and data assimilation components, striving for direct forecasting from observations as initial conditions (vaughan2024aardvark; mcnally2024data).
+However, to date, the forecast-assimilation cycle remains the most reliable mainstream paradigm for operational end-to-end weather forecasting.
+This work primarily concentrates on data assimilation and end-to-end weather forecasting based on conventional observations (e.g., in-situ weather stations, ships and buoys, radiosonde observations), which serve as the most direct and accurate information sources.
+Nowadays, satellite observations are playing an increasingly vital role in data assimilation.
+Typically, they do not directly provide observations of geophysical variables but rather measurements such as radiation, brightness temperature, reflectivity, etc., which require additional inversion steps, although neural networks inherently suit this process (xu2024fuxi).
+Compared with the pre-satellite era, the operational forecasting capabilities of meteorological centers using a complete global observing system including satellites have improved dramatically.
+Lacking the assimilation for satellite and radar observations, especially satellite data that are important for large-scale processes, constitutes a primary limitation of this study and a key direction for our future exploration.
+Another limitation lies in that our method is based on grid data of a given resolution, and interpolating the original off-grid data will inevitably introduce additional errors.
+Recently, techniques based on continuous spatial modeling have offered a promising train of thought to directly handle off-grid data and have achieved impressive results in scaling to arbitrary resolutions (chen2024fnp).
+Moreover, this method of modeling implicit functional representation can support multi-modal data (e.g., in-situ observations and satellite data) to form a universal framework and can be readily extended to 4D data assimilation.
+In addition, since our method does not change the forecasting model, the end-to-end forecasts of the system still inevitably inherit its shortcomings, such as the blurred prediction.
+This problem is hopeful to be solved as the forecasting model improves.
+In fact, ERA5 itself integrates various observational data, including conventional observations, satellites, and radars, in the process of being generated.
+Unlike the optimization process of traditional data assimilation methods, supervised learning makes this information inevitably part of the implicit priors introduced in model training.
+For the current end-to-end paradigm, the precision of the supervised labels determines the upper limit of the method’s performance, posing a challenge to be addressed in the future.
+Producing high-quality analysis data and forecasts that may be superior to any existing product should be the ultimate goal of data assimilation and data-driven revolution, which still requires a lot of exploration and efforts.
+Recent study has shown promising progress and may become the mainstream paradigm for the next stage of the research in the domain of end-to-end weather forecasting (alexe2024graphdop).
+The reanalysis datasets offer the highest-quality complete labels at present and hold an irreplaceable significance.
+In fact, high-quality data has always been essential for the development of AI and many other technologies.
+Therefore, the development of physical models and traditional algorithms remains crucial.
+The purpose of our study is not to replace the traditional NWP systems but to demonstrate the potential of AI methods in tackling real-world challenges.
+Both physical techniques and AI methods have their own strengths and limitations and should be treated equally and considered as powerful tools for benefiting science and society.
 We believe that this is a meaningful exploration and look forward to the deployment and implementation of AI methods in operational systems in the future.
@@
 ERA5 is a global atmospheric reanalysis archive containing hourly weather variables such as temperature, geopotential, wind speed, humidity, etc.
-A subset of ECMWF’s ERA5 dataset for 40 years, from 1979 to 2018, is chosen to train and evaluate the model.
-We choose to conduct experiments on a total of 69 variables at a resolution of 0.25° ($721\times 1440$ grid points), including five upper-air variables with 13 pressure levels (i.e., 50hPa, 100hPa, 150hPa, 200hPa, 250hPa, 300hPa, 400hPa, 500hPa, 600hPa, 700hPa, 850hPa, 925hPa, and 1000hPa), and four surface variables.
-Specifically, the upper-air variables are geopotential (z), temperature (t), specific humidity (q), zonal component of wind (u) and meridional component of wind (v), whose 13 sub-variables at different vertical level are presented by abbreviating their short name and pressure levels (e.g., z500 denotes the geopotential at a pressure level of 500 hPa), and the surface variables are 10-meter zonal component of wind (u10), 10-meter meridional component of wind (v10), 2-meter temperature (t2m) and mean sea level pressure (msl).
-Following a common protocol in the simulation experiments, the data from 1979-2015 are used for training, 2016-2017 for validation, and 2018 for testing.
-The real observational data is parsed from the prepbufr files of GDAS, which is archived by National Centers for Environmental Information (NCEI).
-GDAS provides various types of observational data starting from 2012-02-13 00:00UTC, including multi-source data such as surface land reports, surface marine reports, satellite-derived wind reports, etc.
-Each datum contains the longitude and latitude coordinates, pressure level, time, and the value and quality marker of state variables.
-Regarding the large number of missing GDAS observations after 2018, we use the data from 2017 for testing and others for training to demonstrate the quality of the analysis for the entire year.
+A subset of the ERA5 dataset for 39 years, from 1979 to 2017, is chosen to train and evaluate the model.
+We choose to conduct experiments on a total of 69 variables at a resolution of 0.25° ($721\times 1440$ grid points), including 5 upper-air variables with 13 pressure levels (i.e., 50hPa, 100hPa, 150hPa, 200hPa, 250hPa, 300hPa, 400hPa, 500hPa, 600hPa, 700hPa, 850hPa, 925hPa and 1000hPa), and 4 surface variables.
+Specifically, the upper-air variables are geopotential ($z$), temperature ($t$), specific humidity ($q$), zonal component of wind ($u$) and meridional component of wind ($v$), whose 13 sub-variables at different vertical level are presented by abbreviating their short name and pressure levels (e.g., $z500$ denotes the geopotential at a pressure level of 500 hPa), and the surface variables are 10-meter zonal component of wind ($u10$), 10-meter meridional component of wind ($v10$), 2-meter temperature ($t2m$) and mean sea level pressure (mslmsl).
+We use the data from 1979-2015 for training, 2016 for validation, and 2017 for testing.
+The real observational data are parsed from the prepbufr files of GDAS, which are archived by the National Centers for Environmental Information (NCEI).
+GDAS provides various types of observational data starting from 2012-02-13 00:00UTC, including multi-source data such as aircraft and upper-air reports, surface land/marine reports, scatterometer data, satellite-derived wind reports, etc.
+The prepbufr files consist of many messages with different types, each of which contains many observation columns with different height levels.
+The conventional types of observations in GDAS are used for data assimilation and evaluation in our experiments, covering the main geophysical variables such as temperature, humidity, and wind speed.
+Each observation contains the longitude and latitude coordinates, pressure level, time, and the value and quality marker of the geophysical variables.
+Before being input into the neural network, the observations are first processed into 0.25° grid data through nearest neighbor interpolation, and the grid values without observations are filled with 0.
+The confidence matrix is specified based on the quality marker, time offset, and interpolation distance to represent the quality of the observation value, and the confidence at places without observations is set to 0 to indicate filled spurious observation values.
+Regarding the large number of missing GDAS observations after 2018, data from 2017 are used for testing and others for training.
 Evaluation metrics
-The latitude-weighted root mean square error (RMSE) and the latitude-weighted anomaly correlation coefficient (ACC) are statistical metrics widely used in geospatial analysis and atmospheric science, which are both computed against ground truth data.
-And latitude weighting is a common strategy to account for the varying area represented by different latitudes on a spherical Earth.
-RMSE evaluates the accuracy of a model’s forecasts or estimates of variables across different latitudes, and ACC reflects the model’s ability to capture anomalies (departures from the long-term averaged climatology).
-Regarding the bias drift of AI models, the latitude-weighted mean error (Bias) is also used to assess the bias of estimates against the ground truth [17], and the Bias of an ideal system should be close to zero.
-The definitions of the metrics are as below.
-RMSE
+The latitude-weighted root mean square error (RMSE) is a statistical metric widely used in geospatial analysis and atmospheric science, which evaluates the accuracy of a model’s forecasts or estimates of variables across different latitudes.
+The latitude weighting is a common strategy to account for the varying area represented by different latitudes on a spherical Earth.
+The definition of RMSE is as follows.
 Given the estimate $\hat{x}_{h,w,c}^{t}$ and its ground truth $x_{h,w,c}^{t}$ for the $c$-th channel at time $t$, the RMSE is defined as
@@
 $\operatorname{RMSE}(c,\tau)=\frac{1}{T}\sum_{i=1}^{T}\sqrt{\frac{1}{H_{0}\cdot W_{0}}\sum\nolimits_{h,w}H_{0}\cdot\frac{\operatorname{cos}(\alpha_{h,w})}{\sum_{h^{\prime}=1}^{H_{0}}\operatorname{cos}(\alpha_{h^{\prime},w})}(x_{h,w,c}^{i+\tau}-\hat{x}_{h,w,c}^{i+\tau})^{2}}$
-where $T$ is the total number of test time slots.
-ACC
-Given the forecast $\hat{x}_{h,w,c}^{i+\tau}$ and its ground truth $x_{h,w,c}^{i+\tau}$, and the climatological mean over the day-of-year $C_{h,w,c}^{i+\tau}$ for the $c$-th channel with lead time $\tau$, the ACC is defined as
-$\operatorname{ACC}(c,\tau)=\frac{1}{T}\sum_{i=1}^{T}\frac{\sum\nolimits_{h,w}H_{0}\cdot\frac{\operatorname{cos}(\alpha_{h,w})}{\sum_{h^{\prime}=1}^{H_{0}}\operatorname{cos}(\alpha_{h^{\prime},w})}(x_{h,w,c}^{i+\tau}-C_{h,w,c}^{i+\tau})(\hat{x}_{h,w,c}^{i+\tau}-C_{h,w,c}^{i+\tau})}{\sqrt{\sum\nolimits_{h,w}H_{0}\cdot\frac{\operatorname{cos}(\alpha_{h,w})}{\sum_{h^{\prime}=1}^{H_{0}}\operatorname{cos}(\alpha_{h^{\prime},w})}(x_{h,w,c}^{i+\tau}-C_{h,w,c}^{i+\tau})^{2}\sum\nolimits_{h,w}H_{0}\cdot\frac{\operatorname{cos}(\alpha_{h,w})}{\sum_{h^{\prime}=1}^{H_{0}}\operatorname{cos}(\alpha_{h^{\prime},w})}(\hat{x}_{h,w,c}^{i+\tau}-C_{h,w,c}^{i+\tau})^{2}}}$
-where $C_{h,w,c}^{i+\tau}$ is averaged from 1993 to 2016 with ERA5 data [13, 14].
-Bias
-Given the estimate $\hat{x}_{h,w,c}^{t}$ and its ground truth $x_{h,w,c}^{t}$ for the $c$-th channel at time $t$, the Bias is defined as
-$\operatorname{Bias}(c,t)=\frac{1}{H_{0}\cdot W_{0}}\sum\nolimits_{h,w}H_{0}\cdot\frac{\operatorname{cos}(\alpha_{h,w})}{\sum_{h^{\prime}=1}^{H_{0}}\operatorname{cos}(\alpha_{h^{\prime},w})}(x_{h,w,c}^{t}-\hat{x}_{h,w,c}^{t})$
+where $T$ is the total number of test time slots. The RMSE calculated at the stations is not latitude-weighted.
 Deep network modules
@@
 The patch embedding and patch recovery are implemented through standard convolution and transposed convolution, respectively.
-The patch size of Adas in our experiments is set to 6 and the number of vertical layers remain unchanged.
-This means a patch has $1\times 6\times 6$ pixels for 3D upper-air variables and $6\times 6$ pixels for 2D surface variables.
+The patch size of Adas in our experiments is set to 6 and the number of vertical layers remains unchanged.
+This means that a patch has $1\times 6\times 6$ pixels for 3D upper-air variables and $6\times 6$ pixels for 2D surface variables.
 The kernel sizes and strides of 3D and 2D convolution are the same as the patch size, and the two parts are then concatenated together in the vertical direction.
@@
 Patch merging and patch expanding
-In order to capture multi-scale meteorological features, the patch merging and patch expanding modules are used for down-sampling and up-sampling (Fig.
-2d).
-The patch merging firstly rearrange pixels to reduce the horizontal dimensions by half, while the number of channels increases.
+In order to capture multi-scale meteorological features, the patch merging and patch expanding modules are used for down-sampling and up-sampling (Figure 2d).
+Patch merging first rearranges pixels to reduce the horizontal dimensions by half, while the number of channels increases.
 After the regularization layer, the number of channels is then transformed by a linear layer.
-Therefore, the patch merging achieves halving of the horizontal dimensions and doubling of the channel dimension.
-The patch expanding module performs the opposite operation and makes the encoder form a symmetrical Unet [45] structure.
+Therefore, patch merging achieves halving of the horizontal dimensions and doubling of the channel dimension.
+The patch expanding module performs the opposite operation and makes the encoder form a symmetric Unet (ronneberger2015u) structure.
 Similarly, the number of vertical layers is not changed in the process of patch merging and patch expanding.
 Gated convolution
-Gated convolution is a convoluional operation based on soft-gating, which employs continuous masks ranging between 0 and 1 to represent the degree of validity for each pixel.
-Originally, the gating mask in gated convolution was learned automatically from data at each layer, and the output features were weighted by it.
+Gated convolution (yu2019free) is a convoluional operation based on soft-gating, which employs continuous masks ranging between 0 and 1 to represent the degree of validity for each pixel.
+Originally, the gating mask in gated convolution was learned automatically from the data at each layer, and the output features were weighted by it.
 In Adas, the confidence matrix of our method can serve as the gating mask.
@@
 where $\odot$ represents element-wise multiplication.
-The $Sigmoid(\cdot)$ activation ensures that the values of the confidence matrix are between 0 and 1, and the activation funtion for features can be arbitrary.
-We use the $SiLU(\cdot)$ activation here, which has been proven to be a better choice than $ReLU(\cdot)$ [46, 47].
+The $Sigmoid(\cdot)$ activation ensures that the confidence matrix values are between 0 and 1, and the activation function for the features can be arbitrary.
+We use the $SiLU(\cdot)$ activation here, which has been proven to be a better choice than $ReLU(\cdot)$ (hendrycks2016gaussian).
 The gated convolution module can provide a more complete representation for observations and improve the quality of feature fusion.
@@
 After calculating the cross-attention, the remaining proportion of the background that does not participate in the operation will be added back.
-Conversely, the background fully influences observations since all values in the background have relative high confidence levels generally, which can correct the observations with low confidence levels.
-Therefore, the confidence matrix needs to be updated accordingly, which is also implemented through a convolutional layer with $Sigmoid(\cdot)$ activation (Fig.
-2c).
+In contrast, the background fully influences observations since all values in the background generally have relatively high confidence levels, which can correct the observations with low confidence levels.
+Therefore, the confidence matrix needs to be updated accordingly, which is also implemented through a convolutional layer with $Sigmoid(\cdot)$ activation (Figure 2c).
 The process of gated cross-attention can be expressed as
@@
 The mean absolute error (MAE) loss, also known as L1 loss, is employed to supervise the training of the neural network.
-In multivariate optimization, an issue of imbalanced optimization arises when there are significant differences in the magnitudes of losses for different variables.
+In multi-variate optimization, an issue of imbalanced optimization arises when there are significant differences in the magnitudes of losses for different variables.
 Therefore, it is necessary to weigh the losses for different variables.
 Instead of setting the weights manually, the losses for each variable (i.e., each channel) are automatically weighted to have the same magnitude in our training.
-The models in simulation experiments are trained for 50 epochs using the AdamW optimizer [48] and OneCycleLR scheduler [49].
-The learning rate starts from 1e-6, warms to a maximum value of 1e-4 for ten epochs, and then decays gradually to 1e-10.
+The models in simulation experiments are trained for 50 epochs using the AdamW optimizer (loshchilov2018fixing) and OneCycleLR scheduler (smith2019super).
+The learning rate starts from 1e-6, warms to a maximum value of 1e-4 for 10 epochs, and then decays gradually to 1e-10.
 To ensure adequate training with available GDAS observational data, the models in GDAS experiments are fine-tuned for 50 epochs based on the transfer learning strategy.
@@
 In our system, Adas requires only about 0.6 seconds for the inference of data assimilation on an NVIDIA Tesla-A100 GPU, and FengWu needs about 1.5 seconds for inferring a single-step forecast.
-That is to say, our system can complete a periodic iteration in seconds once the observational data are available, which achieves significant efficiency improvement (about 1000-10000 times faster) compared to traditional data assimilation algorithms [50].
-[…]
+This means that our system can complete a periodic iteration in seconds once the observational data are available, which achieves significant efficiency improvement (about 1000-10000 times faster) compared to traditional data assimilation algorithms (zhang2019operational).
 Acknowledgements
-This research is supported by National Natural Science Foundation of China (No.
-62071127, and 62101137), National Key Research and Development Program of China (No.
-2022ZD0160100), Shanghai Natural Science Foundation (No.
-23ZR1402900) and Shanghai Municipal Foundation (No.2021SHZDZX0103).
-We acknowledge the ECMWF and NCEI for their great efforts in storing and providing the invaluable data, which is very important for this work and the research community.
-We would also like to express our appreciation to Mr.
+This research is supported by National Natural Science Foundation of China (No.62071127 and 62101137), National Key Research and Development Program of China (No.2022ZD0160101), Shanghai Natural Science Foundation (No.23ZR1402900), Shanghai Science and Technology Commission Explorer Program Project (24TS1401300) and Shanghai Municipal Science and Technology Major Project (No.2021SHZDZX0103).
+We acknowledge the ECMWF and NCEI for their great efforts in storing and providing the invaluable data, which are very important for this work and the research community.
+We would also like to express our appreciation to Dr.
+Ben Fei, Mr.
 Huihang Sun and Prof.
-Jiayuan Fan for their suggestion and valuable discussions during the conduction of this research, and the research team and service team in the Shanghai Artificial Intelligence Laboratory for the provision of computational resources and infrastructure.
-Declarations
-[…]
-Competing interests The authors declare no competing interests.
-[…]
-Data availability The ERA5 dataset can be downloaded from the official website of Climate Data Store (CDS) at https://cds.climate.copernicus.eu/.
+Jiayuan Fan for their suggestions and valuable discussions during the conduct of this research, and the research team and service team in the Shanghai Artificial Intelligence Laboratory for the provision of computational resources and infrastructure.
+Data availability
+The ERA5 dataset can be downloaded from the official website of Climate Data Store (CDS) at https://cds.climate.copernicus.eu.
 The GDAS observational bufr files can be obtained from the Archive Information Request System (AIRS) of NCEI at https://www.ncei.noaa.gov/has/HAS.DsSelect.
-[…]
-Code availability The neural network model is developed using standard libraries in open-source platforms including PyTorch.
-Codes and model checkpoints used in this study will be available at https://github.com/kunc3301/FengWu-Adas if ready.
-[…]
-Author contribution L.B.
-and T.C.
-are co-corresponding authors.
-L.B., T.C., J.-J.L.
-and W.O.
-supervised this research.
-Kun Chen, L.B., F.L.
-and P.Y.
-conceived and initiated this project.
-Kun Chen and L.B.
-designed the details of model and developed the methodologies.
-Kun Chen and Kang Chen developed the Python code.
-Kun Chen, Y.X.
-and T.H.
-prepared and processed the data.
-Kun Chen performed experiments.
-Kun Chen, L.B., F.L., P.Y.
-and H.C.
-wrote the manuscript.
-All authors contributed to analysis of data, discussions about the results, and improvement of the presentation.
-    Appendix A Extended Data
-    Figure 7: RMSE and Bias skill of FengWu-Adas with random masks at each time step in simulation experiments.
-Rows 1 and 3 show the RMSE of the analyses for main variables over the year, and rows 2 and 4 show the Bias for them.
-There are no obvious performance differences compared to Fig.
-3, which prove the assimilation ability and robustness of Adas.
-See the caption of Fig.
-3 for all other details.
-    Figure 8: Average RMSE skill of FengWu-Adas for all variables in simulation experiments.
-The different colors represent the RMSE of analyses with different observation ratios, as marked in the legend.
-The first subplot includes four surface variables, and other subplots correspond to five upper-air variables, as indicated in the subplot titles.
-The x-axis represents the single variable (and pressure level).
-The y-axis represents the RMSE.
-    Figure 9: RMSE skill of FengWu-Adas with random initialization in simulation experiments.
-The different colors represent the results with different observation ratios, as marked in the bottom.
-Each subplot represents a single variable (and pressure level), as indicated in the subplot titles, and the subgraph provides a clearer display of the first ten days.
-The x-axis represents the operation time at 6-hour steps over the year.
-The y-axis represents the RMSE, which uses a coordinate axis with non equidistant logarithmic form to ensure visual perception.
-The results further demonstrate the ability of Adas to propagate sparse observation information and the stability of the system.
-    Figure 10: RMSE and ACC skill of FengWu-Adas for weather forecasting over ten days in simulation experiments.
-The different colors represent the forecasts whose initial states are generated after assimilating observations with different ratios, as marked in the bottom.
-Based on the analysis produced by Adas, FengWu can still generate accurate and consistent global medium-range weather forecasts.
-See the caption of Fig.
-6 for all other details.
-    Figure 11: Average RMSE skill of FengWu-Adas for all variables in GDAS experiments.
-The different colors represent the RMSE of analyses after assimilating all data and non-satellite data, as marked in the legend.
-See the caption of Extended Data Fig.
-8 for all other details.
+The IGRA dataset and its documentation are publicly available at https://www.ncei.noaa.gov/data/integrated-global-radiosonde-archive.
+Code availability
+The neural network model is developed using standard libraries in open-source platforms including PyTorch.
+Codes and model checkpoints used in this study are available at https://github.com/OpenEarthLab/FengWu-Adas.